# Occlusion-Aware Seamless Segmentation

### v1

@@ Abstract
Panoramic images can broaden the Field of View (FoV), occlusion-aware prediction can deepen the understanding of the scene, and domain adaptation can transfer across viewing domains. In this work, we introduce a novel task, Occlusion-Aware Seamless Segmentation (OASS), which simultaneously tackles all these three challenges. For benchmarking OASS, we establish a new human-annotated dataset for Blending Panoramic Amodal Seamless Segmentation, i.e., BlendPASS. Besides, we propose the first solution UnmaskFormer, aiming at unmasking the narrow FoV, occlusions, and domain gaps all at once. Specifically, UnmaskFormer includes the crucial designs of Unmasking Attention (UA) and Amodal-oriented Mix (AoMix). Our method achieves state-of-the-art performance on the BlendPASS dataset, reaching a remarkable mAPQ of and mIoU of . On public panoramic semantic segmentation datasets, i.e., SynPASS and DensePASS, our method outperforms previous methods and obtains and in mIoU, respectively. The fresh BlendPASS dataset and our source code will be made publicly available at OASS.
@@
Previous approaches [94, 98, 5, 61, 55] are proposed to address the above challenges separately, resulting in sub-optimal and non-seamless solutions. For example, as the comparison depicted in Fig. 2, methods [94, 98] proposed to address panoramic semantic segmentation cannot handle the object occlusion, whereas methods [61, 55, 5] proposed for amodal segmentation cannot generalize well to panoramic imagery and the FoV occlusion remains unsolved. To address the challenges seamlessly, we propose a novel unmasking transformer framework called UnmaskFormer. The novelty is three-fold: (1) An Unmasking Attention (UA) constructed by a self-attention and an enhanced pooling layer for occlusion prediction, is proposed to unmask the object occlusions within the whole UnmaskFormer framework. (2) We delve deeper into the design of the Deformable Patch Embedding (DPE) [94] and find an alternative yet better solution for addressing the image distortion of panoramas at different stages within the transformer model. (3) We propose an Amodal-oriented Mix (AoMix) method that aims to seamlessly integrate the pinhole and panoramic domains, addressing the challenges posed by inconsistent cross-domain scenes. This method also enhances the model’s capacity to reconstruct invisible regions of occluded objects, ultimately allowing it to unmask the occlusion of perspective within a scene. Based on these crucial designs, our UnmaskFormer can better handle different panorama-based scene understanding tasks, especially solving object occlusion and image distortion. As shown in Fig. 1(b), our UnmaskFormer manages to attain striking performances across Semantic (mIoU), Instance (mAP), Amodal Instance (mAAP), Panoptic (mPQ), and Amodal Panoptic Segmentation (mAPQ) tasks via one model.
In the literature, there is a lack of datasets or benchmarks suitable for learning and evaluating Occlusion-Aware Seamless Segmentation (OASS) methods. To facilitate this evaluation, we spent a large effort to collect and manually annotate a dataset for Blending Panoramic Amodal Seamless Segmentation (BlendPASS). There are panoramic images with 360° FoV and a resolution, captured in street scenes from cities worldwide, featuring both driving and walking perspectives. BlendPASS facilitates the learning optimization of segmentation adaptation in an unsupervised fashion, which unfolds as efficient for dense prediction in panoramic imagery. To establish the evaluation benchmark for the OASS task, panoramic images have been manually annotated precisely at the pixel level.
@@
We introduce a new panorama-based segmentation task, i.e., Occlusion-Aware Seamless Segmentation (OASS), aiming at unmasking the narrow field of view, unmasking the occlusion of perspective, and unmasking the gap of domain in a seamless manner.
@@ 2 Related Work
Domain adaptive panoramic segmentation. Segmentation on wide-FoV fisheye images [13, 66, 85, 63, 86] and panoramic images [23, 43, 80, 83, 100] enables holistic understanding of 360° surroundings [79]. To address the scarcity of annotations, researchers have revisited wide-FoV segmentation from the perspective of unsupervised domain adaptation [38, 32, 84, 82, 67] by leveraging rich training sources from conventional narrow-FoV data. A variety of adaptation methods including self-training [28, 29, 103, 27, 75, 91] and adversarial training [53, 51] methods are studied for panoramic segmentation. In this line, P2PDA [93] employs attention-regulated uncertainty-aware adaptation. Trans4PASS [94, 95] introduces distortion-aware transformers and mutual prototypical adaptation with SAM [40]. DPPASS [99] presents a dual-path solution to overcome style and format gaps. DATR [98] captures neighboring distributions without any geometric constraints. Moreover, panoramic panoptic segmentation [18, 34, 54] is investigated by using techniques like contrastive learning for harvesting generalization benefits [33]. In this work, we aim to enable occlusion-informed understanding with both FoV-wise and spatially amodal reasoning of urban scenes.
Amodal scene segmentation. Amodal instance segmentation is a derivative task of instance segmentation [25, 12, 4, 68, 71, 46, 16], to predict visible regions of objects and their occluded regions simultaneously. Li et al. [42] introduce the concept of amodal instance segmentation, achieved through iterative regression. Zhu et al. [102] derive two amodal segmentation datasets from BSDS [3] and COCO [48]. Qi et al. [61] derive the amodal version of KITTI [21], introducing an independent occlusion classification branch. SLN [97] incorporates a semantics-aware distance map to implement amodal segmentation. ORCNN [17] calculates visible masks to infer occlusion masks. Further, shape and contour priors are frequently leveraged for refining amodal segmentation [19, 6, 45, 76, 44]. Aside from amodal instance segmentation, amodal semantic understanding [102, 31, 5] is an important direction. Along this line, [55, 56] propose a fusion of semantic and amodal instance segmentation. Sekkat et al. [64] uses the CARLA simulator [14] to create a virtual multi-task amodal perception dataset. Different from these works, we intend to achieve occlusion-aware seamless segmentation, which breaks the occlusion limits in terms of both field-of-view and in-field object occlusions.
@@ 3.1 Overview of the Benchmark
In this work, we establish a novel benchmark specifically designed for Occlusion-Aware Seamless Segmentation (OASS). In particular, we address OASS from the perspective of UDA by adapting from the label-rich pinhole domain to the label-scarce panoramic domain. 360° panoramas have a wide FoV and many small objects in the images, which exaggerate the costs of creating pixel-wise annotations in unconstrained surroundings. Our benchmark overcomes the scarcity of amodal segmentation testbeds in panoramic imagery, while addressing three aforementioned challenges: how to address the severe distortions of panoramas when unmasking the narrow field of view, how to predict the full segmentation of objects when unmasking the occlusion of perspective, and how to facilitate optimization of segmentation adaptation when unmasking the gap of domain.
@@ 3.2 BlendPASS
We introduce a novel dataset for Blending Panoramic Amodal Seamless Segmentation (BlendPASS) tailored for OASS. BlendPASS comprises an unlabeled training set of panoramic images for optimizing domain adaptation and a labeled test set of panoramic images. These images are captured from panoramic cameras in driving scenes across cities for the training set and cities for the test set, all at a resolution of pixels. As depicted in Fig. 3(a), we have provided pixel-wise annotations for five distinct visual segmentation tasks, which greatly extends the semantic labels from DensePASS [53]. These annotations cover categories that align with the Cityscapes [11] and are further categorized into Stuff (road, sidewalk, building, wall, fence, pole, light, sign, vegetation, terrain, and sky) and Thing (person, rider, car, truck, bus, train, motorcycle, and bicycle). To ensure the precision of annotations and the rational handling of occluded object parts, we meticulously conduct manual annotation for the test set, with three annotators following the “independent annotating cross-checking voting” workflow, as outlined in [55]. Finally, as illustrated in Fig. 3(b), objects are annotated in the Thing class, with of these objects exhibiting occlusion.
@@ 4 Methodology
In this section, we describe the proposed method UnmaskFormer tailored for OASS, as illustrated in Fig. 5. We start by introducing the principal architecture of UnmaskFormer, featuring Unmasking Attention (UA) blocks crafted for distortion- and occlusion-aware reasoning in OASS. Subsequently, we elaborate on our cross-domain adaptation approach, leveraging the Amodal-oriented Mix (AoMix) method to enhance occlusion-aware knowledge transfer.
@@ 4.1 UnmaskFormer
Apart from the distortion-aware design, we design the UA block by combining the self-attention layer and the enhanced pooling layer. As shown in Fig. 5, after the self-attention layer, the feature is forwarded to calculate the pooling query , the key and the value . After having the self-attended pooling feature , a sigmoid function is applied to calculate the occlusion-aware mask . Then, the pool-attended occlusion-aware mask is used to perform a dot product operation with feature to obtain the occlusion-aware feature . After the UA block, the occlusion-aware feature is further forwarded to an MLP layer as self-attention blocks [77, 94].
@@ 4.2 Cross-Domain Adaptation
UDA strategy. In the OASS setting, the model is assigned to adapt from a labeled source domain to an unlabeled target domain for multiple segmentation tasks, which are images from pinhole and panoramic cameras respectively. To address the domain gap arising from dissimilar data distributions between the source pinhole and target panoramic domains, a comprehensive loss is utilized for training the overall network :
where and represent the source domain supervision loss and the target domain adaptation loss, respectively. is utilized for supervision to learn fundamental segmentation capacity through labeled source samples , whereas the leverages unlabeled target images to enhance segmentation in the target panoramic scenario. includes semantic loss, instance loss, and amodal instance loss, each originating from respective branches. The semantic branch employs cross-entropy loss to assign each pixel to its respective category. As for the instance and amodal instance branches, we follow the conventional approach from Mask R-CNN [26], employing supervision using proposals-based bounding boxes and instance-level masks.
@@
The parameters of the teacher model are updated using the parameters of the student model by the exponential moving average at each iteration :
@@
where denoted as a step function. For a source image , we sum the corresponding amodal instance masks to obtain a binary mask of Thing region. Using the random binary mask and the Thing region binary mask , we fill the source image to reconstruct the masked source image .
Amodal-oriented masked image modeling aims to enhance the model’s ability to reconstruct object regions obscured by realistic object shapes, enabling it to learn information about invisible parts in the scene. To adapt the model to unlabeled target scenes, we randomly sample half of the semantic classes from a source image and paste the corresponding pixels of the masked source image onto a target image based on the source semantic map, creating a new masked mixed image . This approach effectively transfers the model’s capability tailored for distortion- and occlusion-aware to the target panoramic domain.
@@ 5.1 Experiment Setups
Following DAFomer [27], we train UnmaskFomer using the AdamW optimizer [50]. The learning rate is set to , with a weight decay of and linear warmup for iterations followed by polynomial decay. The model is trained on a batch size of with a crop size of for iterations. The EMA decay is set as . We use the same data augmentation parameters as DACS [72] and set the RCS temperature to to enhance sample frequency for rare classes of the source domain. Moreover, we adopt the ImageNet feature distance loss from DAFormer with a weight of . For instance and amodal instance branches, we set all loss weights to . Our experiments are conducted on an NVIDIA Tesla V100 GPU and implemented using PyTorch.
@@
5.2 Results of OASS on KITTI360BlendPASS
We deliver the OASS results on the KITTI360-APSBlendPASS benchmark in Tab. 1. Representative UDA panoptic segmentation methods UniDAPS [96] and EDAPS [62] with our amodal extensions are benchmarked. Further, state-of-the-art panoramic semantic segmentation models Trans4PASS [94] and DATR [98] with distortion-adaptive capacities are compared. It is worth noting that UniDAPS does not support occlusion-aware reasoning with a contour-based decoder.
As shown in Tab. 1, compared with the best-performing UDA panoptic segmentation model EDAPS [62], UnmaskFormer outstrips it by clear margins of in mIoU and in mAPQ. Compared to the panoramic segmentation model Trans4PASS [94], our method also yields significant gains. Delving into the OASS results in semantic- and amodal panoptic segmentation, all benchmarked methods suffer from accurately predicting the full segmentation in the cases considering occlusions, in particular on small objects, evidenced by the unsatisfactory scores in class-wise APQ. Yet, UnmaskFormer harvests great improvements in contrast to the Source-Only model, e.g., on safety-critical pedestrians, truck, and two-wheeler. Finally, under the challenging occlusion scenarios in unstructured surroundings, UnmaskFormer clearly stands out and leads to state-of-the-art OASS scores of in mIoU and in mAPQ. In Fig. 6, we visualize OASS results produced by our UnmaskFormer and state-of-the-art methods DATR [98], Trans4PASS [94], EDAPS [62]. Compared to them, UnmaskFormer excels in segmenting occluded objects and realistically reconstructing vehicle shapes thanks to our UA-based backbone and AoMix design.
@@ 5.3 Results of Panoramic Semantic Segmentation
SynPASS dataset. We further investigate the generalization capacity of the proposed UnmaskFormer on panoramic semantic segmentation. In Tab. 4, we conduct a comparison between known convolutional and transformer models for panoramic semantic segmentation on the SynPASS benchmark [95]. Here, the models are learned on the training set of images and evaluated on the validation set of images with diverse weather and illumination conditions. It can be seen that UnmaskFormer equipped with the unmasking attention and the semantic head, produces state-of-the-art performance measured in mIoU, which reaches . Moreover, in each adverse condition, UnmaskFormer yields robust segmentation of the full panoramas on SynPASS.
@@
DensePASS dataset. As shown in Tab. 4, we further compare UnmaskFormer against efficient panoramic segmentation transformers including DPPASS [99], DATR [98], and Trans4PASS [95], where the models are trained on Cityscapes [11] and tested on the original panoramic semantic segmentation benchmark DensePASS [53]. Compared to the state-of-the-art DATR [98] constituted by distortion-aware attention with parameters, our UnmaskFormer with fewer parameters () attains higher segmentation performance. Our UnmaskFormer achieves in mIoU, which outperforms its counterparts and maintains as a lightweight segmenter. Therefore, UnmaskFormer is suitable for efficiency-relevant panoramic driving scene understanding.

### v2

@@ Abstract
Panoramic images can broaden the Field of View (FoV), occlusion-aware prediction can deepen the understanding of the scene, and domain adaptation can transfer across viewing domains. In this work, we introduce a novel task, Occlusion-Aware Seamless Segmentation (OASS), which simultaneously tackles all these three challenges. For benchmarking OASS, we establish a new human-annotated dataset for Blending Panoramic Amodal Seamless Segmentation, i.e., BlendPASS. Besides, we propose the first solution UnmaskFormer, aiming at unmasking the narrow FoV, occlusions, and domain gaps all at once. Specifically, UnmaskFormer includes the crucial designs of Unmasking Attention (UA) and Amodal-oriented Mix (AoMix). Our method achieves state-of-the-art performance on the BlendPASS dataset, reaching a remarkable mAPQ of and mIoU of . On public panoramic semantic segmentation datasets, i.e., SynPASS and DensePASS, our method outperforms previous methods and obtains and in mIoU, respectively. The fresh BlendPASS dataset and our source code are available at https://github.com/yihong-97/OASS.
@@
Previous approaches [89, 93, 4, 56, 50] are proposed to address the above challenges separately, resulting in sub-optimal and non-seamless solutions. For example, as the comparison depicted in Fig. 2, methods [89, 93] proposed to address panoramic semantic segmentation cannot handle the object occlusion, whereas methods [56, 50, 4] proposed for amodal segmentation cannot generalize well to panoramic imagery and the FoV occlusion remains unsolved. To address the challenges seamlessly, we propose a novel unmasking transformer framework called UnmaskFormer. The novelty is three-fold: (1) An Unmasking Attention (UA) constructed by a self-attention and an enhanced pooling layer for occlusion prediction, is proposed to unmask the object occlusions within the whole UnmaskFormer framework. (2) We delve deeper into the design of the Deformable Patch Embedding (DPE) [89] and find an alternative yet better solution for addressing the image distortion of panoramas at different stages within the transformer model. (3) We propose an Amodal-oriented Mix (AoMix) method that aims to seamlessly integrate the pinhole and panoramic domains, addressing the challenges posed by inconsistent cross-domain scenes. This method also enhances the model’s capacity to reconstruct invisible regions of occluded objects, ultimately allowing it to unmask the occlusion of perspective within a scene. Based on these crucial designs, our UnmaskFormer can better handle different panorama-based scene understanding tasks, especially solving object occlusion and image distortion. Fig. 1(b) shows that UnmaskFormer has striking performances across Semantic (mIoU), Instance (mAP), Amodal Instance (mAAP), Panoptic (mPQ), and Amodal Panoptic Segmentation (mAPQ) tasks by using one model.
To facilitate this evaluation, we spent a large effort to collect and manually annotate a dataset for Blending Panoramic Amodal Seamless Segmentation (BlendPASS). There are panoramic images with 360° FoV and a resolution, captured in street scenes. BlendPASS facilitates the learning optimization of segmentation adaptation in an unsupervised fashion, which unfolds as efficient for dense prediction in panoramic imagery. To establish the evaluation benchmark for the OASS task, panoramic images have been manually annotated precisely at the pixel level.
@@
We introduce a new panorama-based segmentation task, i.e., OASS, aiming at unmasking the narrow field of view, unmasking the occlusion of perspective, and unmasking the gap of domain in a seamless manner.
@@ 2 Related Work
Domain adaptive panoramic segmentation. Segmentation on wide-FoV fisheye images [11, 61, 80, 58, 81] and panoramic images [20, 39, 75, 78, 95] enables holistic understanding of 360° surroundings [74]. To address the scarcity of annotations, researchers have revisited wide-FoV segmentation from the perspective of UDA [34, 28, 79, 77, 62] by leveraging rich training sources from conventional narrow-FoV data. A variety of adaptation methods including self-training [24, 25, 98, 23, 70, 86] and adversarial training [48, 47] methods are studied for panoramic segmentation. In this line, P2PDA [88] employs attention-regulated uncertainty-aware adaptation. Trans4PASS [89, 90] introduces distortion-aware transformers and mutual prototypical adaptation with SAM [36]. DPPASS [94] presents a dual-path solution to overcome style and format gaps. DATR [93] captures neighboring distributions without any geometric constraints. Moreover, panoramic panoptic segmentation [15, 30, 49] is investigated by using techniques like contrastive learning for harvesting generalization benefits [29]. In this work, we aim to enable occlusion-informed understanding with both FoV-wise and spatially amodal reasoning of urban scenes.
Amodal scene segmentation. Amodal instance segmentation is a derivative task of instance segmentation [21, 10, 3, 63, 66, 42, 13, 45], to predict visible regions of objects and their occluded regions simultaneously. Li et al. [38] introduce the concept of amodal instance segmentation, achieved through iterative regression. Qi et al. [56] derive the amodal version of KITTI [18], introducing an independent occlusion classification branch. SLN [92] incorporates a semantics-aware distance map to implement amodal segmentation. ORCNN [14] calculates visible masks to infer occlusion masks. Further, shape and contour priors are frequently leveraged for refining amodal segmentation [16, 5, 41, 71, 40]. Aside from amodal instance segmentation, amodal semantic understanding [97, 27, 4] is an important direction. Along this line, [50, 51] propose a fusion of semantic and amodal instance segmentation. Sekkat et al. [59] uses the CARLA simulator [12] to create a virtual multi-task amodal perception dataset. Different from these works, we intend to achieve occlusion-aware seamless segmentation, which breaks the occlusion limits in terms of both field-of-view and in-field object occlusions.
@@ 3.1 Overview of the Benchmark
In this work, we establish a novel benchmark specifically designed for OASS. In particular, we address OASS from the perspective of UDA by adapting from the label-rich pinhole domain to the label-scarce panoramic domain. 360° panoramas have a wide FoV and many small objects in the images, which exaggerate the costs of creating pixel-wise annotations in unconstrained surroundings. Our benchmark overcomes the scarcity of amodal segmentation testbeds in panoramic imagery, while addressing three aforementioned challenges: how to address the severe distortions of panoramas when unmasking the narrow field of view, how to predict the full segmentation of objects when unmasking the occlusion of perspective, and how to facilitate optimization of segmentation adaptation when unmasking the gap of domain.
@@ 3.2 BlendPASS
We introduce a novel dataset for Blending Panoramic Amodal Seamless Segmentation (BlendPASS) tailored for OASS. BlendPASS comprises an unlabeled training set of panoramic images for optimizing domain adaptation and a labeled test set of panoramic images. These images are captured from panoramic cameras in driving scenes, all at a resolution of pixels. As depicted in Fig. 3, we have provided pixel-wise annotations for five distinct visual segmentation tasks, which greatly extends the semantic labels from DensePASS [48]. Specifically, we have further annotated the instance and amodal instance labels. These labels cover categories that align with the Cityscapes [9] and are further categorized into Stuff (road, sidewalk, building, wall, fence, pole, light, sign, vegetation, terrain, and sky) and Thing (person, rider, car, truck, bus, train, motorcycle, and bicycle). To ensure the precision of annotations and the rational handling of occluded object parts, we meticulously conduct manual annotation for the test set, with three annotators following a cross-checking process. Finally, objects are annotated in the Thing class, with of these objects exhibiting occlusion. More details can be found in the supplementary.
@@ 4.1 UnmaskFormer
Apart from the distortion-aware design, we design the UA block by combining the self-attention layer and the enhanced pooling layer. As shown in Fig. 5, after the self-attention layer, the feature is forwarded to calculate the pooling query , the key and the value , where and is the global average pooling operator. After having the self-attended pooling feature , a sigmoid function is applied to calculate the occlusion-aware mask . Then, the pool-attended occlusion-aware mask is used to perform a dot product operation with feature to obtain the occlusion-aware feature . After the UA block, the occlusion-aware feature is further forwarded to an MLP layer as self-attention blocks [72, 89].
@@ 4.2 Cross-Domain Adaptation
UDA strategy. In the OASS setting, the model is assigned to adapt from a labeled source domain to an unlabeled target domain for multiple segmentation tasks, which are images from pinhole and panoramic cameras respectively. To address the domain gap arising from dissimilar data distributions between the source pinhole and target panoramic domains, a comprehensive loss is utilized for training the overall network , where and represent the source domain supervision loss and the target domain adaptation loss, respectively. is utilized for supervision to learn fundamental segmentation capacity through labeled source samples , whereas the leverages unlabeled target images to enhance segmentation in the target panoramic scenario. includes semantic loss, instance loss, and amodal instance loss, each originating from respective branches. The semantic branch employs cross-entropy loss to assign each pixel to its respective category. As for the instance and amodal instance branches, we follow the conventional approach from Mask R-CNN [22], employing supervision using proposals-based bounding boxes and instance-level masks.
@@
The parameters of the teacher model are updated using the parameters of the student model by the EMA at each iteration.
@@
where denoted as a step function. For a source image , we sum the corresponding amodal instance masks to obtain a binary mask of Thing region. Using the random binary mask and the Thing region binary mask , we fill the source image to reconstruct the masked source image . Amodal-oriented masked image modeling aims to enhance the model’s ability to reconstruct object regions obscured by realistic object shapes, enabling it to learn information about invisible parts. To adapt to unlabeled target scenes, we randomly sample half of the semantic classes from a source image and paste the corresponding pixels of the masked source image onto a target image based on the source semantic map, creating a new masked mixed image . This approach effectively transfers the model’s capability tailored for distortion- and occlusion-aware to the target panoramic domain.
@@ 5.1 Experiment Setups
Following DAFomer [23], we train UnmaskFomer using AdamW [46]. The learning rate is set to , with a weight decay of and linear warmup for iterations followed by polynomial decay. The model is trained on a batch size of with a crop size of for iterations. More details can be found in the supplementary.
@@
5.2 Results of OASS on KITTI360-APSBlendPASS
We deliver the OASS results on the KITTI360-APSBlendPASS benchmark in Tab. 1. Representative UDA panoptic segmentation methods UniDAPS [91] and EDAPS [57] with our amodal extensions are benchmarked. Further, SOTA panoramic semantic segmentation models Trans4PASS [89] and DATR [93] with distortion-adaptive capacities are compared. It is worth noting that UniDAPS does not support occlusion-aware reasoning with a contour-based decoder.
As shown in Tab. 1, compared with the best-performing UDA panoptic segmentation model EDAPS [57], UnmaskFormer outstrips it by clear margins of in mIoU and in mAPQ. Compared to the panoramic segmentation model Trans4PASS [89], our method also yields significant gains. Delving into the OASS results in semantic- and amodal panoptic segmentation, all benchmarked methods suffer from accurately predicting the full segmentation in the cases considering occlusions, in particular on small objects, evidenced by the unsatisfactory scores in class-wise APQ. Yet, UnmaskFormer harvests great improvements in contrast to the Source-Only model, e.g., on safety-critical pedestrians, truck, and two-wheeler. Finally, under the challenging occlusion scenarios in unstructured surroundings, UnmaskFormer clearly stands out and leads to SOTA OASS scores of in mIoU and in mAPQ. In Fig. 6, we visualize OASS results produced by our UnmaskFormer and SOTA methods DATR [93], Trans4PASS [89], EDAPS [57]. Compared to them, UnmaskFormer excels in segmenting occluded objects and realistically reconstructing vehicle shapes thanks to our UA-based backbone and AoMix design.
@@ 5.3 Results of Panoramic Semantic Segmentation
We further investigate the generalization capacity of the proposed UnmaskFormer using two panoramic semantic segmentation datasets [90, 48].
SynPASS. In Tab. 4, we conduct a comparison between known convolutional and transformer models for panoramic semantic segmentation on the SynPASS [90]. Here, the models are learned on the training set and evaluated on the validation set with diverse weather and illumination conditions. UnmaskFormer equipped with the UA and the semantic head, produces SOTA performance in mIoU, which reaches . Moreover, in each adverse condition, UnmaskFormer yields robust segmentation of the full panoramas on SynPASS.
@@
DensePASS. As shown in Tab. 4, we further compare UnmaskFormer against efficient panoramic segmentation transformers including DPPASS [94], DATR [93], and Trans4PASS [90], where the models are trained on Cityscapes [9] and tested on the original DensePASS [48]. Compared to the DATR [93] constituted by distortion-aware attention with parameters, our UnmaskFormer with fewer parameters () attains higher segmentation performance. Our UnmaskFormer achieves in mIoU, which outperforms its counterparts and maintains as a lightweight segmenter.